# Counterfactuality, back-action, and information gain in multi-path interferometers

### v1

@@ II Statistical analysis of counterfactual effects
In this Section, we show that quantum mechanical counterfactuals differ from classical counterfactual logic due to the loss of coherence between the states that have been excluded by the observation of a negative result, and the states which are not excluded. We then develop a way to analyse the benefit such coherence provides for counterfactual discrimination tasks—the ‘counterfactual gain’.
@@
Consider a scenario where we are trying to guess correctly whether an absorber is present or absent. We are assuming this can be treated as being completely unbiased—based on the information we have, it is equiprobable whether or not the absorber was inserted. This assignment of equiprobability can be justified in a Bayesian sense, as our prior being complete ignorance as to whether or not the absorber was inserted. If our guess is based on a single trial, the optimal strategy is to compare the probability $P(m)$ without the absorber to the probability $P(m|X_{a})$ in the presence of an absorber. If the absorber reduces the probability of the outcome ($P(m|X_{a})<P(m)$), the best guess upon observing a photon arriving at output $m$ is that the absorber was absent. While classically the absorber can only either reduce the probability or keep it the same, to consider quantum scenarios we also need to take into account the possibility that the presence of the absorber somehow increases the probability of the outcome $m$ ($P(m|X_{a})>P(m)$), in which case the best guess upon observing $m$ is that the absorber is present. These counter-intuitive outcomes are essential for a counterfactual gain in our ability to identify the presence of a detector: even though the photon was not absorbed, the outcomes $m$ with $P(m|X_{a})>P(m)$ provide (non-classical) evidence for the presence of the absorber. We can now quantify the total effect of this counterfactual gain by finding the average error probability. Since our guess will be wrong whenever the outcome $m$ was caused by the less likely condition, this error probability is given by
@@
In the case of a classical counterfactual situation ($P(m|X_{a})<P(m)$ for all $m$), the presence of the absorber will not be detected unless the photon is actually absorbed. This results in an average error probability of $(1-P(a))/2$ for a classical counterfactual situation, where the absorption simply removes some of the photons from the outputs $m$. For the evaluation of the counterfactual gain, it is convenient to express the error probability using the statistical distance,
@@
$P_{\text{error}}=\frac{1}{2}-\frac{1}{2}\Delta_{a}$ (3)
@@ IV Back-action
correctly reproducing the original relation in Eq. (5) where the fsubtraction of two times the Kirkwood-Dirac term is split into a term representing the removal of photons in $a$ from the outcomes $m$, while the other half represents part of the back-action effect $\chi_{B}/2$, which ensures that no negative probabilities will be observed. It seems characteristic of quantum counterfactuals that the quasiprobability subtracted to represent the absence of the absorbed photons in the output also appears as part of the modification of the probabilities caused by the unavoidable back-action effect that provides the counterfactual gain.
@@ V Statistics and back-action in Hofmann’s Three-Path Interferometer
Is it possible to distinguish the effects of non-classical statistics from the additional back-action effects by considering the relation between measurement contexts? One might object that the negative value of the Kirkwood-Dirac term $\varrho(F,3)=-1/9$ could be replaced by a different quasiprobability able to explain the counterfactual effects. As shown in the previous section, this is only prevented by the observation that the Elitzur-Vaidman terms represent the propagation from $F$ to the output ports when the other ports (e.g. $P_{2}$ and $S_{2}$) are blocked, and these terms would have to be described by the same quasiprobability in order to explain the counterfactual effects. In the present scenario the Elitzur-Vaidman terms for all outputs are $1/27$, so the prediction for a removal of these particles would be a homogeneous reduction of all output probabilities to $8/27$. Conversely, the distribution of particles from $F$ in the outputs would have to be $(5/27,5/27,-7/27)$ according to the counterfactual effects observed when $F$ is blocked. Consistency between the distributions observed when path $F$ is blocked and when only $F$ is open requires that the back-action re-distributes the particles from low probability outputs to high probability outputs, with the change in probability given by $\chi_{B}=(-2/27,-2/27,4/27)$.

### v2

@@ II Statistical analysis of counterfactual effects
In this Section, we show that quantum mechanical counterfactuals differ from classical counterfactual logic due to the loss of coherence between the states that have been excluded by the observation of a negative result, and the states which are not excluded. This difference can be quantified in terms of a quantum advantage in the task of discriminating between the presence of an absorber and its absence in a specific path of a single-photon interferometer by looking at which output of that interferometer, if any, the photon arrived at. Since this quantum advantage is associated with quantum mechanical counterfactuals, we will refer to it as ‘counterfactual gain’ in the following.
@@
Consider a scenario where we are trying to guess correctly whether an absorber is present or absent. We are assuming this can be treated as being completely unbiased—based on the information we have, it is equiprobable whether or not the absorber was inserted. This assignment of equiprobability can be justified in a Bayesian sense, as our prior being complete ignorance as to whether or not the absorber was inserted. If our guess is based on a single trial, the optimal strategy is to compare the probability $P(m)$ without the absorber to the probability $P(m|X_{a})$ in the presence of an absorber. If the absorber reduces the probability of the outcome ($P(m|X_{a})<P(m)$), the best guess upon observing a photon arriving at output $m$ is that the absorber was absent. While classically the absorber can only either reduce the probability or keep it the same, to consider quantum scenarios we also need to take into account the possibility that the presence of the absorber somehow increases the probability of the outcome $m$ ($P(m|X_{a})>P(m)$), in which case the best guess upon observing $m$ is that the absorber is present. These counter-intuitive outcomes are essential for a counterfactual gain in our ability to identify the presence of a detector: even though the photon was not absorbed, the outcomes $m$ with $P(m|X_{a})>P(m)$ provide (non-classical) evidence for the presence of the absorber. We can now quantify the total effect of this counterfactual gain by finding the average error probability for the optimal guessing strategy described above. Since our guesses will be based on the most likely condition for the respective outcome, wrong guesses will occur whenever the outcome was caused by the less likely condition. If the detector is present, any detection in an outcome $m$ with a probability of $P(m|X_{a})<P(m)$ will result in an error. If the detector is absent, any detection in an outcome $m$ with a probability of $P(m)<P(m|X_{a})$ will result in an error. Assuming that the presence or absence of the detector is equally likely (as justified above), the average error probability is obtained by
@@
In the case of a classical counterfactual situation ($P(m|X_{a})<P(m)$ for all $m$), the presence of the absorber will not be detected unless the photon is actually absorbed. This results in an average error probability of $(1-P(a))/2$ for a classical counterfactual situation, where the absorption simply removes some of the photons from the outputs $m$. For the evaluation of the counterfactual gain, it is convenient to express the error probability using a quantitative expression of the difference between the two probability distributions $\{P(m)\}$ of outcomes without the absorber and $\{P(a),P(m|X_{a})\}$ of outcomes in the presence of the absorber, where $P(a)$ again is the probability of absorption. A commonly used measure of statistical distance is the total variation distance, defined as one half of the sum over the absolute values of the differences between the probabilities of each outcome. In the present case, the outcome $a$ indicates absorption and has a probability of zero in the absence of the absorber. The statistical distance is therefore given by
@@
$P_{\text{error}}=\frac{1}{2}-\frac{1}{2}\Delta_{a}.$ (3)
@@ IV Back-action
correctly reproducing the original relation in Eq. (5) where the subtraction of two times the Kirkwood-Dirac term is split into a term representing the removal of photons in $a$ from the outcomes $m$, while the other half represents part of the back-action effect $\chi_{B}/2$, which ensures that no negative probabilities will be observed. It seems characteristic of quantum counterfactuals that the quasiprobability subtracted to represent the absence of the absorbed photons in the output also appears as part of the modification of the probabilities caused by the unavoidable back-action effect that provides the counterfactual gain.
@@ V Statistics and back-action in Hofmann’s Three-Path Interferometer
Is it possible to distinguish the effects of non-classical statistics from the additional back-action effects by considering the relation between measurement contexts? One might object that the negative value of the Kirkwood-Dirac term $\varrho(F,3)=-1/9$ could be replaced by a different quasiprobability able to explain the counterfactual effects. As shown in the previous section, this is only prevented by the observation that the Elitzur-Vaidman terms represent the propagation from $F$ to the output ports when the other ports (e.g. $P_{2}$ and $S_{2}$) are blocked, and these terms would have to be described by the same quasiprobability in order to explain the counterfactual effects. In the present scenario the Elitzur-Vaidman terms for all outputs are $1/27$, so the prediction for a removal of these particles would be a homogeneous reduction of all output probabilities to $8/27$. Conversely, the distribution of particles from $F$ in the outputs would have to be $(5/27,5/27,-7/27)$ according to the counterfactual effects observed when $F$ is blocked to explain the scenario in this way. This would be nonsensical, as path $F$ would have to somehow distribute -7/27 of a particle to output 3. Consistency between the distributions observed when path $F$ is blocked and when only $F$ is open requires that the back-action re-distributes the particles from low probability outputs to high probability outputs, with the change in probability given by $\chi_{B}=(-2/27,-2/27,4/27)$.